The Optical Scam Company has forecast a sales growth rate of 25 percent for next year. The current financial statements are shown here Income Statement Sales Costs $ 32,000,000 26,309,400 Taxable income Taxes $ 5,690,600 1,991,710 Net income $3,698,890 Dividends Addition to retained earnings $ 1,479,556 2,219,334 Balance Sheet Assets Liabilities and Equity Current assets $ 7,360,000 Short-term debt Long-term debt $7,040,000 2,240,000 Fixed assets 18,560,000 Common stock Accumulated retained earnings $ 6,070,000 10,570,000 $ 16,640,000 $ 25,920,000 Total equity Total assets$25,920,000 Tota liabilities and equity a. Calculate the external financing needed for next year. (Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.) External financing needed b-1. Prepare the firms pro forma balance sheet for next year. (Do not round intermediate calculations and round your answers to the nearest whole number, e.g., 32.) Balance Sheet Assets Current assets Liabilities and equity Short-term debtp0

Long-term debt Fixed assets Common stock Accumulated retained earnings Total equity Total assets Total liabilities and equity b-2. Calculate the external financing needed. (Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.) External financing needed c. Calculate the sustainable growth rate for the company based on the current financial statements. (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g, 32.16.) Sustainable growth rate
Answer #1
a) EFN = [(Assets / Sales) x Δ sales] - [(Debt / Sales) x Δ sales] - [(Projected sales x Profit margin) x (1-Dividend payout ratio)]
= (25920000/32000000)*32000000*25%-(7040000/32000000)*32000000*25%-32000000*125%*11.56%*60% =
= $1945600 (EFN)
b-1) Year Current Year Basis for projections Next year
INCOME STATEMENT:
Sales 32000000 +25% 40000000
Costs 26309400 82.22% of sales 32886750
EBIT 5690600 7113250
Taxes at 35% 1991710 2489638
Net income 3698890 0.115590313 4623613
Dividends (40%) 1479556 1849445
Addition to retained earnings 2219334 2774168
BALANCE SHEET:
Current assets 7360000 23% of sales 9200000
Fixed assets 18560000 58% of sales 23200000
25920000 32400000
Short term debt 7040000 22% of sales 8800000
Long term debt 2240000 2240000
Common stock 6070000 6070000
Accumulated retained earnings 10570000 +2774168 13344168
Total debt and equity 25920000 30454168
b-2) EFN 1945833
Difference in answers under [a] and [b-2] is due to approximation.
c) SGR = ROE*b/(1-ROE*b)
where ROE = return on equity,
b = retention ratio.
ROE = 3698890/(6070000+10570000) = 22.23%
Retention ratio = 60%
SGR = 0.2223*0.60/(1-0.2223*0.60) = 15.39%